OpenStudy (anonymous):

Henry rolls 2 number cubes numbered 1 through 6 while playing his favorite board game. He will get a second turn if he rolls a sum that is an even number less than 10. What are Henry's chances of getting a second turn when he rolls the number cubes? seven over eighteen 11 over 18 5 over 36 17 over 36 I thought it was 4/11???

OpenStudy (smored):

1+1=2 1+3=4 1+5=6 2+2=4 2+6=8 3+3=6 3+5=8 4+4=8 and they're all less than 10

OpenStudy (anonymous):

crap, It's just been a few months since I've done any math, I've got it now. I don't know what i was thingking, thanks!

OpenStudy (smored):

no prob